# AgentClinic: a multimodal agent benchmark to evaluate AI in simulated clinical environments

### v4

@@ F.2 Does the patient language model affect accuracy?
We find that the diagnostic accuracy drops from to 52% with a GPT-4 doctor and GPT-4 patient agent to 48% with a GPT-4 doctor and a GPT-3.5 patient agent. The accuracy with a GPT-4 doctor and Mixtral patient agent is similarly reduced to 46%. Inspecting the dialogues, we noticed that the GPT-3.5 patient agent is more likely to repeat back what the doctor has asked. For example, consider the following dialogue snippet: "Doctor: Have you experienced any muscle twitching or cramps? Patient: No, I haven’t experienced any muscle twitching or cramps." Now consider this dialogue from a GPT-4 patient agent: "Doctor: Have you had any recent infections, like a cold or the flu, before these symptoms started? Patient: Yes, I’ve had a couple of colds back to back and a stomach bug in the last few months." We find that, while GPT-4 also partakes in doctor rehearsal, GPT-4 patient agents are more likely to reveal additional symptomatic information than GPT-3.5 agents which may contribute to the higher accuracy observed with GPT-4-based patient agents.
@@ Appendix I OSCE Examination Structure
Experiential learning in the context of AI agents refers to the ability to accumulate knowledge and insights from past interactions and apply them to future tasks (Wang et al. (2024); Zhao et al. (2024)). This technique allows agents to improve their performance over time by learning from successes, failures, and feedback received during previous engagements. This was previously explored in Agent Hospital (Li et al. (2024)) through an experience retrieval system. By maintaining a form of "memory" or knowledge base that updates through interaction, agents can become better at handling similar situations, adapting to user preferences, and providing increasingly relevant and accurate responses as they gain more "experience" in their operational domain. In our work, we enable the doctor agent to use a memory "notebook" which persists across patients. Here, the doctor agent can write useful tips such as the following example from the doctor agent: "[Note #17] Remember that timing and onset of symptoms can provide valuable diagnostic insights."
@@
To enable the doctor agent to research medical information, we introduce a method using an adaptive form of retrieval augmented generation (RAG) from medical sources. RAG involves retrieving relevant information from a knowledge base and using it to augment the input of an LLM during the generation process (Gao et al. (2023)), thereby improving the factual consistency of generated text by grounding it in retrieved information. Conventional RAG methods passively retrieve information at every inference call without allowing the agent to control the timing or content of retrieval. To address this limitation, we employ adaptive retrieval (Jiang et al. (2023); Asai et al. (2023)), which enables the LLM to actively determine when and what information to retrieve. Our implementation provides the doctor agent with two categories of retrieval: internet and textbook databases. The internet database contains material from sources such as PubMed111https://pubmed.ncbi.nlm.nih.gov/ research articles, StatPearls222https://www.statpearls.com/—a database of articles written for healthcare professionals—and Wikipedia articles on various medical topics. The textbook database includes 18 medical textbooks commonly used by medical students in the United States (Jin et al. (2021)). The doctor agent can retrieve information by issuing commands similar to requesting medical scans, using the format: "Research [database] [search query]". For example, the command "Research textbooks ’What are the symptoms of myasthenia gravis?’" prompts the retrieval of relevant information (see Appendix B.3 for more detail).

### v5

@@ F.2 Does the patient language model affect accuracy?
We find that the diagnostic accuracy drops from to 52% with a GPT-4 doctor and GPT-4 patient agent to 48% with a GPT-4 doctor and a GPT-3.5 patient agent. The accuracy with a GPT-4 doctor and Mixtral patient agent is similarly reduced to 46%. Inspecting the dialogues, we noticed that the GPT-3.5 patient agent is more likely to repeat back what the doctor has asked. For example, consider the following dialogue snippet: “Doctor: Have you experienced any muscle twitching or cramps? Patient: No, I haven’t experienced any muscle twitching or cramps." Now consider this dialogue from a GPT-4 patient agent: “Doctor: Have you had any recent infections, like a cold or the flu, before these symptoms started? Patient: Yes, I’ve had a couple of colds back to back and a stomach bug in the last few months." We find that, while GPT-4 also partakes in doctor rehearsal, GPT-4 patient agents are more likely to reveal additional symptomatic information than GPT-3.5 agents which may contribute to the higher accuracy observed with GPT-4-based patient agents.
@@
F.3 Coverage of MedQA content versus AgentClinic-MedQA
To better understand the performance differences between MedQA and AgentClinic-MedQA, we conducted an analysis to quantify the amount of relevant patient information obtained by the doctor agents in each setting. Specifically, we focused on measuring coverage—the proportion of relevant information successfully extracted by the doctor agent through dialogue with the patient agent or through measurement interactions.
For this analysis, we selected a sample of MedQA cases and their corresponding AgentClinic-MedQA simulations, using GPT-4 as the doctor agent. In MedQA, all relevant patient information, such as symptoms, medical history, and test results, is provided upfront in a static format. In contrast, AgentClinic-MedQA requires the doctor agent to dynamically gather this information through interactions. To evaluate coverage, we manually reviewed the dialogues in AgentClinic-MedQA and determined whether the doctor agent extracted each piece of relevant information identified in the MedQA cases. Coverage was calculated as the ratio of extracted information to the total relevant information available in the MedQA cases.
Our findings revealed that the average coverage in AgentClinic-MedQA was 67%. Furthermore, the coverage was notably higher (72%) in cases where the doctor agent provided a correct diagnosis, compared to 63% in cases where the diagnosis was incorrect. These results suggest that the ability to extract more complete information is a key factor in accurate diagnoses in AgentClinic-MedQA. The discrepancy in diagnostic accuracy between MedQA and AgentClinic-MedQA can likely be attributed to the additional complexity of acquiring information in the latter, as opposed to the static format of the former.
F.4 Multi-Agent Evaluations
To explore the role of multi-agent collaboration in clinical diagnosis, we benchmarked two novel multi-agent frameworks: Multi-Agent Debate (Du et al. (2023)) and MedAgents (Tang et al. (2023)), across three language model configurations: GPT-4, GPT-4o, and Claude-3.5-Sonnet. These frameworks aim to emulate team-based diagnostic settings by incorporating multiple interacting agents, enabling structured collaboration and debate to refine diagnostic outcomes.
Multi-Agent Debate
: This approach allows multiple doctor agents to debate and converge on a diagnosis, leveraging diverse reasoning pathways (Du et al. (2023)). We observe that Claude-3.5-Sonnet achieves the highest diagnostic accuracy with 64.1% ± 3.4, outperforming both GPT-4 (51.7% ± 3.0) and GPT-4o (37.9% ± 3.1). These results highlight Claude-3.5-Sonnet’s collaborative reasoning capabilities, likely attributable to its higher inter-agent consistency and adaptability in resolving conflicting diagnostic opinions.
MedAgents
: This framework promotes collaborative decision-making through structured task delegation among agents, simulating multidisciplinary team interactions in clinical settings (Tang et al. (2023)). Again, Claude-3.5-Sonnet leads with an accuracy of 65.2% ± 3.6, followed by GPT-4 (53.1% ± 3.1) and GPT-4o (40.1% ± 3.3). The improved performance across all configurations compared to single-agent baselines suggests that task specialization among agents enables more comprehensive data collection and interpretation, particularly when supported by robust collaboration mechanisms.
@@ Appendix I OSCE Examination Structure
Experiential learning in the context of AI agents refers to the ability to accumulate knowledge and insights from past interactions and apply them to future tasks (Wang et al. (2024); Zhao et al. (2024)). This technique allows agents to improve their performance over time by learning from successes, failures, and feedback received during previous engagements. This was previously explored in Agent Hospital (Li et al. (2024)) through an experience retrieval system. By maintaining a form of “memory" or knowledge base that updates through interaction, agents can become better at handling similar situations, adapting to user preferences, and providing increasingly relevant and accurate responses as they gain more “experience" in their operational domain. In our work, we enable the doctor agent to use a memory “notebook" which persists across patients. Here, the doctor agent can write useful tips such as the following example from the doctor agent: “[Note #17] Remember that timing and onset of symptoms can provide valuable diagnostic insights."
@@
To enable the doctor agent to research medical information, we introduce a method using an adaptive form of retrieval augmented generation (RAG) from medical sources. RAG involves retrieving relevant information from a knowledge base and using it to augment the input of an LLM during the generation process (Gao et al. (2023)), thereby improving the factual consistency of generated text by grounding it in retrieved information. Conventional RAG methods passively retrieve information at every inference call without allowing the agent to control the timing or content of retrieval. To address this limitation, we employ adaptive retrieval (Jiang et al. (2023); Asai et al. (2023)), which enables the LLM to actively determine when and what information to retrieve. Our implementation provides the doctor agent with two categories of retrieval: internet and textbook databases. The internet database contains material from sources such as PubMed111https://pubmed.ncbi.nlm.nih.gov/ research articles, StatPearls222https://www.statpearls.com/—a database of articles written for healthcare professionals—and Wikipedia articles on various medical topics. The textbook database includes 18 medical textbooks commonly used by medical students in the United States (Jin et al. (2021)). The doctor agent can retrieve information by issuing commands similar to requesting medical scans, using the format: "“Research [database] [search query]". For example, the command “Research textbooks ’What are the symptoms of myasthenia gravis?’" prompts the retrieval of relevant information (see Appendix B.3 for more detail).
@@
| Dataset Name | Sample Size | Modalities Included | Task Types/Descriptions |
| AgentClinic-NEJM | 120 cases derived from NEJM case challenges | Multimodal (Text + Images) | Open-ended diagnostic tasks requiring image analysis and patient dialogues. |
| AgentClinic-MedQA | 215 cases derived from USMLE case challenges | Text | Simulated cases with structured patient information from USMLE data. |
| AgentClinic-MIMIC-IV | 200 cases derived from MIMIC-IV | Text | Simulated cases with structured patient information from real-world EHR data. |
| AgentClinic-Spec | 260 cases derived from from MedMCQA | Text | Specialist diagnostic cases from 9 medical specialties, including pediatrics, psychiatry, and internal medicine. |
| AgentClinic-Lang | 749 cases derived from AgentClinic-MedQA | Multilingual Text | AgentClinic-MedQA cases translated for 7 languages (English, Chinese, Hindi, Korean, Spanish, French, Persian). |
| Model | AgentClinic-MedQA Accuracy (%) |
| Multi-Agent Debate (gpt-4) | 51.7 ± 3.0 |
| Multi-Agent Debate (gpt-4o) | 37.9 ± 3.1 |
| Multi-Agent Debate (claude-3.5-sonnet) | 64.1 ± 3.4 |
| MedAgents (gpt-4) | 53.1 ± 3.1 |
| MedAgents (gpt-4o) | 40.1 ± 3.3 |
| MedAgents (claude-3.5-sonnet) | 65.2 ± 3.6 |